13. The brakes of a 1000 kg automobile can exert a retarding force of 3000N. Find the distance the car will move before stropping if it is traveling at the rate of 24 m/s when the brakes are applied (96 m).


1
Expert's answer
2021-09-25T13:11:08-0400

Given:

"m=1000\;\rm kg"

"F_f=3000\:\rm N"

"v_i=24\:\rm m/s"


The energy-work theorem says

"\frac{mv_f^2}{2}-\frac{mv_i^2}{2}=W=-F_fd"

Hence, the distance traveled by the car to rest

"d=\frac{mv_i^2}{2F_f}=\frac{1000*24^2}{2*3000}=96\:\rm m"